What is the probability of getting a license plate that has a repeated letter or digit if you live in a state that has two numerals followed by five letters followed by one numeral?
Answer by jim_thompson5910(35256)

There are 10*10*26*26*26*26*26*10 = 11,881,376,000 plates possible.

There are 10*9*26*25*24*23*22*8 = 5,683,392,000 plates that have no repeated letters or digits. So there are 11,881,376,000 - 5,683,392,000 = 6,197,984,000 plates that have repeated digits or letters (and are not all unique)

So the probability is 6,197,984,000/11,881,376,000 = 14899/28561 = 0.52165540422254